OpenStudy (jazzyluvsu123):

What is 3 1/2 miles - 3,520 yards ? (Put answer in Miles) MEDAL AND FAN

OpenStudy (anonymous):

1.5 i do believe. 1 mile=1760 yards

OpenStudy (jazzyluvsu123):

How would i work this out tho?

OpenStudy (anonymous):

Mile=5280 feet and Yard=3 feet 5280*3.5=18480 3520*3=10560 18480-10560=7290 7290/5280=1.38

OpenStudy (jazzyluvsu123):

yes i got it ! ty

OpenStudy (anonymous):

No problem, enjoy the rest of your studying. :)

OpenStudy (jazzyluvsu123):

;]

OpenStudy (anonymous):

I did 3.5 x 1760 which i got 6160. Then its just simple subtraction. 6160 - 3520 = 2640. You divide that by a mile 2640 / 1760 which gave me 1.5

OpenStudy (anonymous):

Did i do something wrong?

OpenStudy (anonymous):

No, I messed up my division at the end, probably hit one more number than I had meant to. Both of our methods had worked, and both answer should be the same. Great job on catching my mistake. The answer is 1.5 miles, sorry @JazzyLuvsU123 and @ObeyFCS

OpenStudy (jazzyluvsu123):

Its okay ;p 1.38 in a fractice is 1 1/2 so i got it correct

OpenStudy (jazzyluvsu123):

*fraction

OpenStudy (anonymous):

Alright, sorry again though.

OpenStudy (anonymous):

anytime cx